Andy drove his motorboat downstream for 5 hours before heading back. After cruising back upstream for 7 hours, Andy was 10 miles short of his starting point. If the speed of the river is 5 mph, find the speed of his boat in still water? What distance did Andy cruise downstream before heading back?
Answer by josmiceli(19441)

Let s = speed in still water
(5 + s)*5 = (-5 + s)*7 +10
This equation says that the distance travelled downstream equals
the distance travelled upstream plus 10 miles (rate*time = rate*time + 10 miles)
25 + 5s = -35 + 7s + 10
25 + 35 - 10 = 7s - 5s
50 = 2s
s = 25 miles per hour
The distance travelled downstream is (5 + s)*5
(5 + 25)*5 = 150 miles
check to see that this equals the other side of the equation
150 = (-5 + s)*7 + 10
150 = (-5 + 25)*7 + 10
150 = 140 + 10
150 = 150
OK